ganeshie8 (ganeshie8):

A certain kind of migratory birds start reproducing at the age of \(2\) years. \(2.5\%\) of the baby birds die each year. \(10\%\) of the baby birds that survive become adult each year. The adult birds reproduce at a rate of \(4\%\) per year and die at a rate of \(3\%\) per year. Estimate the number of baby birds and adult birds at the end of \(30\) th year if the starting population of baby birds and adult birds are \(10, 20\) respectively.

OpenStudy (dan815):

system of equations this time, everything else same

OpenStudy (dan815):

2* system

ganeshie8 (ganeshie8):

can these be solved individually by decoupling or something

ganeshie8 (ganeshie8):

anyways lets see if i can setup the recurrence relations

ganeshie8 (ganeshie8):

baby birds \(b_n = b_{n-1}(1-0.025-0.1)\) adult birds \(a_n = a_{n-1}(1+0.04-0.03) + 0.1*b_n\)

OpenStudy (anonymous):

$$\begin{align*}B(n)&=B(n-1)-0.025\cdot B(n-1)+0.04\cdot A(n-1)\\&=0.975\cdot B(n-1)+0.04\cdot A(n-1)\\A(n)&=A(n-1)-0.03\cdot A(n-1)+0.1\cdot B(n-1)\\&=0.97\cdot A(n-1)+0.1\cdot B(n-1)\end{align*}$$ so: $$\begin{bmatrix}B(n)\\A(n)\end{bmatrix}=\begin{bmatrix}0.975&0.04\\0.97&0.1\end{bmatrix}\begin{bmatrix}B(n-1)\\A(n-1)\end{bmatrix}$$ so it follows $$\begin{bmatrix}B(n)\\A(n)\end{bmatrix}=\begin{bmatrix}0.975&0.04\\0.97&0.1\end{bmatrix}^n\begin{bmatrix}B(0)\\A(0)\end{bmatrix}$$

OpenStudy (anonymous):

oops, the bottom row hsould be \(0.1\quad0.97\)

ganeshie8 (ganeshie8):

i think il need to use matlab for this

OpenStudy (anonymous):

hmm, well what we have is not diagonalizable unfortunately so there's no way to decouple this cleanly

ganeshie8 (ganeshie8):

Oh it cannot have independent eigenvectors is it

ganeshie8 (ganeshie8):

but since all the numbers are less than 1, is it okay to assume the matrix approaches 0 for large n ?

OpenStudy (campbell_st):

are they golden plovers.. ..?

OpenStudy (anonymous):

\[ b_n = (1-0.025)(0.9)b_{n-1}+(0.04)a_{n-1}\\ a_n = (1-0.025)(0.1)b_{n-1}+(1-0.03)a_{n-1} \\ A_n = \begin{bmatrix} 0.8775&0.04\\ 0.0975&0.97 \end{bmatrix}^{n-1}A_{1} \]

ganeshie8 (ganeshie8):

Haha similar kind id guess but the saddest thing is that they gonna extinct sooner based on our matrix!

OpenStudy (campbell_st):

ok... blame global warming

OpenStudy (empty):

Plug it into matlab or octave ;P

OpenStudy (anonymous):

hmm, interesting premultiplication of the 0.1 maturation rate @wio. is that matrix diagonalizable

OpenStudy (anonymous):

hmm, well that's no fun

OpenStudy (empty):

I think there's something elegant about being able to use computers :P

ganeshie8 (ganeshie8):

i think wolframalpha is lying, two independent eigenvectors means we can diagonalize http://www.wolframalpha.com/input/?i=eigenvectors+%7B%7B0.8775%2C+0.04%7D%2C%7B0.0975%2C+0.97%7D%7D

OpenStudy (anonymous):

well, the point of asking whether we can diagonalize is because it'll make the matrix power very, very easy :-) and yes, the fact the eigenvalues are below 0 tells us that in the long run the behavior is for the populations to die out

OpenStudy (anonymous):

but if they're not both below 0, then that doesn't necessarily hold